# Data Quality Issues in Flare Prediction using Machine Learning Models

### v1

@@ Abstract
Machine learning models for forecasting solar flares have been trained and tested using a variety of data sources, such as Space Weather Prediction Center (SWPC) operational and science-quality data. Typically, data from these sources is minimally processed before being used to train and validate a forecasting model. However, predictive performance can be impaired if defects in and inconsistencies between these data sources are ignored. For a number of commonly used data sources, together with softwares that query and then output processed data, we identify their respective defects and inconsistencies, quantify their extent, and show how they can affect the predictions produced by data-driven machine learning forecasting models. We also outline procedures for fixing these issues or at least mitigating their impacts. Finally, based on our thorough comparisons of the impacts of data sources on the trained forecasting model in terms of predictive skill scores, we offer recommendations for the use of different data products in operational forecasting.
@@ 1 Introduction
In recent years, machine learning methods have been applied with great success to a variety of space weather forecasting problems, such as solar eruption forecasting (Leka:2018b; leka2019comparison; Camporeale:2019a; chen2019iden; kasapis2022interpretable; whitman2023review; chen2025decent), geomagnetic index prediction (iong2022new; iong2024sparse; chen2025geodgp), and the prediction of terrestrial impacts of space weather events (wang2023forecast; sun2025matrix). These methods have also been used for solar flare forecasting, for which they have improved the state of the art (jiao2020solar; sun2021improved; sun2022predicting; sun2023tensor; chen2024solar). A machine learning method predicts the value of a response variable from the values of one or more predictor variables. In solar flare forecasting in particular, the response is typically the flare class or maximum soft X-ray flux during a flare; the predictors may include remove sensing images such as those from the Solar Dynamics Observatory (SDO) (Boerner:2012a; lemen2012atmospheric; pesnell2012; woods2012; sun2023tensor) or extracted features like the Space eather HMI Active Region Patch (SHARP) parameters (SHARPs). The method is trained to generate predictions via minimization of a loss function that measures the degree to which predicted outcomes deviate from observed outcomes; during training, the mapping from predictors to predictions is manipulated (e.g., by gradient descent) to decrease the loss. Common choices of loss function in the flare forecasting literature include the cross-entropy loss (for binary classification of flares as strong or weak, e.g., chen2019iden), sum of squares loss (for prediction of the logarithm of the peak flux, e.g., jiao2020solar), and the negative log-likelihood (in either classification or regression tasks with statistical models, e.g., sun2023tensor). State-of-the-art binary classification methods for forecasting flares include discriminant analysis (leka2018nwra), logistic regression, support vector machines (bobra2015sola), random forests, gradient boosting machines, long-short term memory (LSTM) neural networks (liu2019predicting), and other more advanced network structures. The literature shows that the performance of machine learning methods is more constrained by the train/validation/test design (random split versus chronological split, for example) and sample size (collecting non-overlapping flare events versus rolling window definition of maximum flaring events) than by the particular machine learning method adopted, due to solar flare data sparsity; see Wang2020lstm; sun2022predicting; chen2024solar for more discussion. Therefore, in this paper, instead of exhaustively evaluating all available machine learning methods, we select representative ones (to be detailed later) and focus on using them to calibrate the impact of the data sources (from different databases) and quality (scientific or operational) we use for training the machine learning model.
@@
Different pairs of predictor and outcome types have been used in solar flare forecasting literature. In Figure 1, we provide an overview of how predictions are produced by a machine learning method from different kinds of predictor data for both classification and regression in solar flare forecasting. Predictors are commonly computed from either images of active regions or summary statistics on those regions. The upper left part of the schematic depicts predictors being computed from Helioseismic and Magnetic Imager (HMI) and (Atmospheric Imaging Assembly) AIA images, while the lower left part depicts predictors being computed from SHARP parameters, which are summary statistics calculated from HMI images. If the method performs classification, then the outcome is the binary variable of whether a flare will occur; if the method performs regression, then the outcome is a number, like the future peak soft X-ray flux. The upper and lower right panels show comparisons of actual outcomes to mock predictions for classification and regression, respectively. The images and plots display data from around the time of an X8.7-class flare that occurred on 14 May 2024 in NOAA active region 13664 (HARP 11149). Here we briefly discuss some relevant works corresponding to the parts illustrated in Figure 1. Detailed numerical results will be presented in Sections 2 and 4.
Image Predictors, Binary Response
Image predictors typically consist of multi-channel AIA and HMI images of the Sun. Full-disk images of the Sun at their native resolution of $4096\times 4096$ for each channel are usually too computationally and memory-intensive to be used as inputs to machine learning models, even when performing batch learning. However, resizing these images through down-sampling or other interpolation methods can degrade spatial perception and erase meaningful, fine-grained details that are useful for flare prediction.
@@
On the other hand, multi-channel images focused on active regions [Jin et al., 2024, in prep] must not only deal with potential concurrent temporal effects (for active regions whose timelines overlap) but also with varying image resolutions, as the sizes and shapes of active regions differ. To address this, works such as Pandey2024 perform a rectangular crop that maximizes unsigned flux (a SHARP parameter). However, other work, such as sun2023tensor, uses a similar but different approach by performing a crop centered on the polarity inversion line (PIL). PILs are typically the main source of flare activity, as shown in many works e.g. Zirin:1993; Green:2018; sun2021improved, so such crops can not only keep predictive power high but also allow for succinct comparisons of features (such as topological and spatial features) the model detects with the derived PIL for potential interpretability results.
Using these multi-channel images of active regions is quite common for flare prediction. In the most basic case, we can develop and train machine learning models to predict whether or not a flare will be “strong” within some time horizon. In this case, the response is binary, with “strong” typically meaning C-class or higher (C+) or M-class or higher (M+). While this kind of forecasting may seem simpler than the kinds described below, it can still be difficult. For example, if only M+ flares are considered strong, forecasting can be difficult because of class imbalance and because of the lack of a clear distinction between flares just below and just above the M-class boundary. Instead of just predicting whether a flare will be strong, we can predict the flare class itself. However, this exacerbates the boundary issue, as it now occurs at multiple boundaries.
@@
Many works use image predictors to predict binary flare strength or flare classes. We briefly review a subset of these. Focusing on works that use only 2D magnetogram imaging data, Park2018 trains a deep Convolutional Neural Network (CNN) on full-disk SDO/HMI magnetograms for binary predictions of C+ flares. Meanwhile, huang2018deep and zheng2019solar train CNNs to do multi-class prediction of C-, M-, and X-class flares using full-disk magnetograms. On the other hand, li2022knowledge and sun2022predicting train CNNs on resized HMI magnetograms focused on active regions. Beyond using only 2D magnetogram data, jonas2018flare uses Gabor filters and other techniques on a subset of the AIA imaging channels to train a linear model to predict M+ binary flare outcomes. Additionally, nishizuka2018deep trains a deep neural network on AIA imaging channels to predict the binary responses for both C+ and M+ flares. For works combining both AIA and HMI images, nishizuka2021operational trains a CNN for binary predictions using both C+ and M+ decision rules on full-disk HMI and AIA images, and sun2023tensor proposes a tensor Gaussian process with contraction to combine AIA and HMI images to predict strong flares (M-class) versus weak flares (B-class).
Image Predictors, Continuous Response
@@
Instead of predicting flare classes, another generalization for flare prediction is to directly predict the peak X-ray flux intensity of a flare, now treating the task as a regression problem rather than classification. There is not as much work done along this line as regression is inherently more difficult than classification, but we still provide a brief review here. In muranushi2015ufcorin, the authors use 2D line-of-sight full-disk magnetogram data and then perform a wavelet transform to train a time-series regression model to predict the peak flux intensity of flares. On the other hand, sun2023tensor uses both AIA and HMI imaging data aligned on active regions to perform a tensor regression to predict peak flux intensity. In these works, the models predict a continuous response, but the predictions are then converted into “classification” outputs using flare-class cutoffs in order to use commonly adopted classification metrics such as accuracy and skill scores. However, a work with which they do not bin their continuous predictions is in vandersande2025thes, which investigates key challenges and best practices that should be kept when treating flare prediction as strictly a regression problem.
Vector Predictors, Binary Response
@@
In addition to high-resolution images that pose challenges to model training, many flare-forecasting studies rely on input data vectors that are derived from the summary statistics or data-driven features computed from the images. Space-Weather HMI Active Region Patches (SHARPs)(SHARPs), derived from photospheric magnetic field data taken by the HMI aboard the SDO, provide 25 magnetic and geometric summary parameters. SHARPs are widely used across flare, SEP, and CME forecasting applications. Similar to SHARPs, Space-Weather MDI Active Region Patches (SMARPs)(SMARPs) are derived from maps of the solar surface magnetic field taken by the Michelson Doppler Imager (MDI) aboard the Solar and Heliospheric Observatory (SOHO).
@@
The binary classification of strong flares (such as $\geq$ M-class) against weak flares (such as $\leq C$ or $\leq B$ class) with multi-dimensional time series predictors is the most prevalent in the data-driven solar flare forecasting literature. Previous work on machine learning and deep learning models trained on multidimensional predictors extracted from photospheric magnetic field data (such as SHARPs) shows that carefully engineered vector physical summaries can already achieve competitive performance (bobra2015sola; florios2018forecasting; liu2019predicting; chen2019iden; nishizuka2018deep). Vector predictors are not limited to features derived from images. (verma2026optimal) studies the probability of a strong flare event as an extreme-event prediction using soft X-ray flux time-series vectors.
Vector Predictors, Continuous Response
@@
Flares are classified based on their peak intensity, which is defined to be the peak value of the GOES X-ray flux during the flare; for example, M-class flares have a peak intensity that is at least ${10}^{-5}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$ and less than ${10}^{-4}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$, and X-class flares have a peak intensity that is at least ${10}^{-4}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$ (goes8-15_readme). A typical formulation of the flare forecasting problem is predicting whether an M+ flare will occur in a future 24-hour period (leka2018sola). In this formulation, the outcome is binary. Table 1 provides a justification for an alternative formulation in which the outcome is continuous. The table is based on the SWAN-SF benchmark dataset, which contains various kinds of data, including flare and GOES X-ray flux data, for the period 2010-2018 (angryk2020multivariate). For each date between 1 May 2010 and 29 December 2018, we checked whether an M+ flare occurred and whether the maximum X-ray flux was at or above the M-class threshold of ${10}^{-5}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$. All but four dates either had both an M+ flare and an M+ maximum X-ray flux or had neither. This shows that predicting whether an M+ flare will occur during a 24-hour window is essentially equivalent to predicting whether the maximum X-ray flux during the window will be above the M-class threshold. Predicting the latter event can be done by forecasting the maximum X-ray flux, which is a continuous outcome.
@@
Table 1: The joint distribution of the M+ maximum flux and flare indicators for dates between 1 May 2010 and 29 December 2018, inclusive. The SWAN-SF benchmark dataset described in angryk2020multivariate was used to compute the counts.
Forecasting a continuous outcome has several advantages over forecasting a binary outcome. First, a prediction of a binary outcome is a prediction of whether an event of interest will happen, but not of how severe the event will be if it does happen. A prediction of a continuous outcome is a prediction of both occurrence and severity. Second, the flare class thresholds are arbitrary; there is no substantive difference between the strongest flares in one class and the weakest flares in the next higher class. If ignored, the lack of a clear distinction between adjacent flare classes can make it difficult for a binary forecasting method to perform well, because the method may have trouble correctly predicting flares near a class boundary.
Several works use a continuous outcome instead of a binary outcome. For example, vandersande2025thes models the maximum soft X-ray flux in a 24-hour window using linear regression, random forests, multilayer perceptrons (MLPs), and convolutional neural networks (CNNs). boucheron2015pred employs support vector regression (SVR) trained on magnetic features extracted from MDI images, using a linearly spaced continuous-valued label vector to represent flare size. jiao2020solar develops a mixed LSTM model to forecast the maximum flare intensity, and considers classification models built on top of the regression.
In the sequel, we will describe issues that need to be addressed when using the various kinds of data mentioned previously and suggest ways to address those issues. In Section 2, we discuss response data, covering various data sources, quality, and a systematic comparison. In Section 3, we discuss predictor data, with image-based predictors being discussed in Section 3.1 and vector predictors being discussed in Section 3.2. Finally, in Section 4, we show how predictive performance can be degraded if the issues we raise aren’t addressed.

### v2

# Defects and Inconsistencies in Solar Flare Data Sources: Implications for Machine Learning Forecasting

@@ Abstract
Machine learning models for forecasting solar flares have been trained and evaluated using a variety of data sources, including Space Weather Prediction Center (SWPC) operational and science-quality data. Typically, data from these sources are minimally processed before being used to train and validate a forecasting model. However, predictive performance can be affected if defects and inconsistencies between these data sources are ignored. For a number of commonly used data sources, together with software that queries and outputs processed data, we identify their respective defects and inconsistencies, quantify their extent, and show how they can affect the predictions produced by data-driven machine-learning forecasting models. We also outline procedures for fixing these issues or at least mitigating their impacts. Finally, based on thorough comparisons of the effects of data sources on the trained forecasting model’s predictive skill scores, we offer recommendations for the use of different data products in operational forecasting.
@@ 1 Introduction
Machine learning methods have been applied with great success to a variety of space weather forecasting problems, such as solar eruption forecasting (e.g. Leka and Barnes, 2018b; Leka et al., 2019; Chen et al., 2019a; Kasapis et al., 2022; Whitman et al., 2023; Chen et al., 2025a), geomagnetic index prediction (e.g. Iong et al., 2022; Hu et al., 2023; Nair et al., 2023; Iong et al., 2024; Chen et al., 2025b), and the prediction of terrestrial impacts of space weather events (e.g. Wang et al., 2019; Telloni et al., 2023; Wang et al., 2023; Elfiky et al., 2025; Sun et al., 2025); see Camporeale et al. (2018) and Camporeale (2019) for reviews of methods and applications. In particular, machine learning methods have been used for solar flare forecasting, showing superior performance as compared with state-of-the-art physics-based models (e.g. Florios et al., 2018; Liu et al., 2019b; Sun et al., 2021; Georgoulis et al., 2021; Nishizuka et al., 2021; Sun et al., 2022; Zheng et al., 2023; Pandey et al., 2023).
@@
A machine learning forecasting method typically predicts the value of a response variable from one or more predictor variables. In solar flare forecasting, in particular, the response is typically the flare class (Nishizuka et al., 2018) or maximum soft X-ray flux during a flare (Yi et al., 2020); the predictors may include remote sensing images such as those from the Solar Dynamics Observatory (SDO) (Boerner et al., 2012; Lemen et al., 2012; Pesnell et al., 2012b; Woods et al., 2012; Sun et al., 2023) or extracted features like the Space eather HMI Active Region Patch (SHARP) parameters (Bobra et al., 2014). The method is trained to generate predictions by minimizing a loss function that quantifies the extent to which predicted outcomes deviate from observed outcomes. During training, the mapping from predictors to predictions is optimized (e.g., via gradient-based methods) to minimize the loss. Common choices of loss function in the flare forecasting literature include the cross-entropy loss (for binary classification of flares as strong or weak, e.g., Chen et al. (2019a)), sum of squares loss (for prediction of the logarithm of the peak flux, e.g., Jiao et al. (2020b)), and the negative log-likelihood (in either classification or regression tasks with statistical models, e.g., Sun et al. (2023)). State-of-the-art binary classification methods for forecasting flares include discriminant analysis (Leka et al., 2018), logistic regression (Korsós et al., 2021), support vector machines (Bobra and Couvidat, 2015), random forests (Liu et al., 2017), gradient-boosting machines (Cinto et al., 2020), long short-term memory (LSTM) neural networks (Liu et al., 2019b), and other more advanced network architectures (e.g. Zheng et al., 2019; Sun et al., 2022; Pandey et al., 2024).
The literature shows that the performance of machine learning methods is more constrained by the train, validation, and test design (random split versus chronological split, for example) and sample size (collecting non-overlapping flare events versus rolling window definition of maximum flaring events) than by the particular machine learning method adopted, due to solar flare data sparsity, especially strong M/X class flares; see Wang et al. (2020b); Sun et al. (2022); Chen et al. (2024) for more discussions. Therefore, in this paper, instead of exhaustively evaluating all available machine learning methods, we select representative ones (to be detailed later) and focus on using them to calibrate the impact of the data sources (from different databases) and quality (scientific or operational) we use for training the machine learning model.
@@
1.1 Predictor-Outcome Combinations in Solar Flare Forecasting
@@
Different pairs of predictor and outcome types have been used in the solar flare forecasting literature. In Figure 1, we provide an overview of how predictions are produced by a machine learning method from different types of predictor data for both classification and regression in solar flare forecasting. Predictors are commonly computed from either the images of active regions or summary statistics on those regions. The upper left part of the schematic depicts predictors being computed from Helioseismic and Magnetic Imager (HMI) and (Atmospheric Imaging Assembly) AIA images, while the lower left part depicts predictors being computed from SHARP parameters, which are summary statistics calculated from HMI measurements. If the method performs classification, the outcome is a binary variable indicating whether a flare will occur; if the method performs regression, the outcome is a numeric variable, such as the future peak soft X-ray flux. The upper and lower right panels show comparisons of actual outcomes to mock predictions for classification and regression, respectively. The images and plots present data from around the time of an X8.7-class flare on 14 May 2024 in NOAA active region 13664 (HARP 11149). Here we briefly discuss some relevant works corresponding to the parts illustrated in Figure 1. Detailed numerical results will be presented in Sections 2 and 4. Here, we briefly review the different predictor-outcome combinations in solar flare forecasting literature with machine learning methods.
1.1.1 Image Predictors and Binary Responses
@@
Image predictors typically consist of multi-channel AIA and HMI images of the Sun. Full-disk images of the Sun at their native resolution of $4096\times 4096$ for each channel are usually too computationally and memory-intensive to serve as inputs to machine learning models, even when using batch learning. However, resizing these images through down-sampling or other interpolation methods can degrade spatial perception and erase meaningful, fine-grained details that are useful for flare prediction. On the other hand, multi-channel images focused on active regions [Jin et al., 2026, in prep] must not only account for potential concurrent temporal effects (i.e., when active regions share timelines) but also for varying image resolutions, as active regions vary in size and shape. To address this, works such as Pandey et al. (2024) perform a rectangular crop that maximizes unsigned flux (a SHARP parameter). However, other work, such as Sun et al. (2023), employs a similar yet distinct approach by performing a crop centered on the polarity inversion line (PIL, Schrijver (2007)). PILs are typically the main source of flare activity, as shown in many works e.g., Zirin and Wang (1993); Green et al. (2018); Sun et al. (2021), so such crops can not only keep predictive power high but also allow for succinct comparisons of features (such as topological and spatial features) that the model detects with the derived PIL for potential interpretability results.
@@
Multi-channel images of active regions are commonly adopted in flare prediction. Typically, machine learning models are trained to forecast whether a flare will be “strong” within a set time period: a binary prediction with response being $\geq M$ (or $\geq C$) flares. Despite its simplicity, binary classification still presents challenges such as class imbalance and ill-defined boundaries between flare categories, particularly near the chosen threshold (e.g., just below or above the M-class). Alternatively, models can try to predict the exact flare class, but this introduces similar boundary problems across multiple class thresholds. Many studies use image predictors to predict binary flare classes. We briefly review a subset of these. Focusing on works that use only 2D magnetogram imaging data, Park et al. (2018) trains a deep Convolutional Neural Network (CNN) on full-disk SDO/HMI magnetograms for binary predictions of C+ flares. Meanwhile, Huang et al. (2018) and Zheng et al. (2019) train CNNs to perform multi-class predictions of C-, M-, and X-class flares using full-disk magnetograms. On the other hand, Li et al. (2022) and Sun et al. (2022) train CNNs on resized HMI magnetograms focused on active regions. Beyond using only 2D magnetogram data, Jonas et al. (2018) applies Gabor filters and other techniques to a subset of AIA imaging channels to train a linear model to predict $\geq M$ binary flare outcomes. Furthermore, Nishizuka et al. (2018) trains a deep neural network on AIA imaging channels to predict the binary responses for both $\geq C$ and $\geq M$ class flares. For works combining both the AIA and HMI images, Nishizuka et al. (2021) trains a CNN for binary predictions using both $\geq C$ and $\geq M$ decision rules on full-disk HMI and AIA images, and Sun et al. (2023) proposes a tensor Gaussian process regression with contraction operators to combine AIA and HMI images to predict strong flares ($\geq$ M-class) versus weak flares ($\leq$ B-class).
1.1.2 Image Predictors and Continuous Responses
@@
Instead of predicting flare classes, another generalization of flare prediction is to directly predict the peak X-ray flux of a flare, treating the task as a regression problem rather than a classification problem. There is less work in this area because regression is inherently more difficult than classification, but we still provide a brief review here. In Muranushi et al. (2015), the authors use 2D line-of-sight full-disk magnetogram data and then perform a wavelet transform to train a time-series regression model to predict the peak flux intensity of flares. On the other hand, Sun et al. (2023) uses both AIA and HMI imaging data aligned on active regions to perform a tensor regression to predict peak flux intensity. In these works, the models predict a continuous response, but the predictions are then converted into “classification” outputs using flare-class cutoffs to enable the use of commonly adopted classification metrics such as accuracy and skill scores. However, a work in which they do not bin their continuous predictions is van der Sande and Muñoz-Jaramillo (2025), which investigates key challenges and best practices to consider when treating flare prediction as a strictly regression problem.
1.1.3 Vector Predictors and Binary Responses
@@
In addition to high-resolution images that pose challenges for model training, many flare-forecasting studies rely on input data vectors derived from summary statistics or data-driven features computed from the images. Space-Weather HMI Active Region Patches (SHARPs)(Bobra et al., 2014), derived from photospheric magnetic field data taken by the HMI aboard the SDO, provide 25 magnetic and geometric summary parameters. SHARPs are widely used for flare forecasting and have recently demonstrated success in SEP predictions (Yu et al., 2025). Similar to SHARPs, Space-Weather MDI Active Region Patches (SMARPs)(Bobra et al., 2021) are derived from maps of the solar surface magnetic field taken by the Michelson Doppler Imager (MDI) aboard the Solar and Heliospheric Observatory (SOHO). Sun et al. (2022) adopts the SMARPs for flare forecasting.
@@
The binary classification of strong flares (such as $\geq$ M-class) against weak flares (such as $\leq C$ or $\leq B$ class) with multi-dimensional time series predictors is the most prevalent in the data-driven solar flare forecasting literature. Previous work on machine learning and deep learning models trained on multidimensional predictors extracted from photospheric magnetic field data (such as SHARPs) shows that carefully engineered vector physical summaries can already achieve competitive performance (Bobra and Couvidat, 2015; Florios et al., 2018; Liu et al., 2019b; Chen et al., 2019a; Nishizuka et al., 2018; Wang et al., 2020b; Nishizuka et al., 2021). Vector predictors are not limited to features derived from images. (Verma et al., 2026) studies the probability of a strong flare event as an extreme-event prediction using soft X-ray flux time-series vectors.
1.1.4 Vector Predictors and Continuous Responses
@@
Flares are classified based on their peak intensity, which is defined to be the peak value of the GOES X-ray flux during the flare. For example, M-class flares have a peak intensity that is at least ${10}^{-5}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$ and less than ${10}^{-4}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$, and X-class flares have a peak intensity that is at least ${10}^{-4}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$ (Mothersbaugh III et al., 2023). A typical formulation of the flare forecasting problem is predicting whether an M+ flare will occur in a future 24-hour period (Leka and Barnes, 2018a). In this formulation, the outcome is binary. Table 1 provides a justification for an alternative formulation in which the outcome is continuous. The table is based on the Space Weather Analytics for Solar Flares (SWAN-SF) benchmark dataset, which contains various kinds of data, including flare and GOES X-ray flux data, for the period 2010-2018 (Angryk et al., 2020). For each date between May 1, 2010 and December 29, 2018, we checked whether an M+ flare occurred and whether the maximum X-ray flux was at or above the M-class threshold of ${10}^{-5}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$. All but four dates either had both an M+ flare and an M+ maximum X-ray flux or had neither. This shows that predicting whether an M+ flare will occur during a 24-hour window is essentially equivalent to predicting whether the maximum X-ray flux during the window will be above the M-class threshold. Predicting the latter event can be achieved by forecasting the maximum X-ray flux, which is a continuous outcome. See Verma et al. (2026) for more discussions.
@@
Table 1: The joint distribution of the M+ maximum flux and flare indicators for dates between May 1, 2010 and December 29, 2018, inclusive. The SWAN-SF benchmark dataset described in Angryk et al. (2020) was used to compute the counts.
Forecasting a continuous outcome has several advantages over forecasting a binary outcome. First, a prediction of a binary outcome is a prediction of whether an event of interest will occur, but not of its severity if it does. A prediction of a continuous outcome involves predictions of both occurrence and severity (see e.g., Jiao et al., 2020b; Verma et al., 2026). Second, the flare class thresholds are arbitrary, meaning that there is no substantive difference between the strongest flares in one class and the weakest flares in the next higher class. If ignored, the lack of a clear distinction between adjacent flare classes can make it difficult for a binary forecasting method to perform well because the method may have trouble correctly predicting flares near a class boundary. Several studies use a continuous outcome rather than a binary outcome. For example, van der Sande and Muñoz-Jaramillo (2025) models the maximum soft X-ray flux in a 24-hour window using linear regression, random forests, multilayer perceptrons (MLPs), and convolutional neural networks (CNNs). Boucheron et al. (2015) employs support vector regression (SVR) trained on magnetic features extracted from MDI images, using a linearly spaced continuous-valued label vector to represent flare size. Jiao et al. (2020b) develops a mixed LSTM model to forecast the maximum flare intensity, and considers classification models built on top of the regression.
1.2 Outline
In the sequel, we will describe the issues that arise when using the various types of data mentioned above in Section 1.1 and suggest ways to address them. In Section 2, we focus on response data, covering various data sources, quality, and a systematic comparison. In Section 3, we focus on predictor data, with image-based predictors discussed in Section 3.1 and vector predictors discussed in Section 3.2. Finally, in Section 4, we show how predictive performance can be degraded if the issues we raise aren’t addressed. We conclude in Section 5 with take-home messages on our numerical comparison results, clarifying strengths and limitations of different data products.